OpenStudy (anonymous):

integrate sen(x).cos^2(x)dx. thanks in advance

OpenStudy (anonymous):

\[-\frac{ 1 }{ 3 }\cos ^3x\]

OpenStudy (anonymous):

i know the answer. what i'm in doubt is about how to reach it. anyway, thank you mhmdrz91.

OpenStudy (kainui):

So for this you might notice that you have the derivative of something multiplied by it, which allows you to do substitution. It will take a little practice to become familiar with it, but after a short while it will be almost obvious. Try using the substitution u=cosx. I'll help you along the way.

OpenStudy (anonymous):

i've already done it. u=\[\cos ^{2} x\], and i got \[du=-2\cos x.\sin x\], \[-du/2= \cos x.\sin x\] it's wrong, isn't it? because you can't reach the right result with this du, i think.

OpenStudy (anonymous):

put cos^x = z diff both sides , -sinxdx=dz put dat in your equation, \[\int\limits_{}^{}z ^{2}dz\] dat equals \[-z ^{3}\div3 + c\] put d value of z ..

OpenStudy (kainui):

I already suggested u=cosx. If one substitution doesn't work, try to figure out why, since it obviously looks very close it might be very similar. So like I said, cosx. It's not cos^2x.

OpenStudy (anonymous):

sorry its only cosx=z and the answer will be -(1/3)cos^3x +c

OpenStudy (kainui):

I think we all 4 of us know what the answer is.

OpenStudy (anonymous):

ya and i solved it .

OpenStudy (kainui):

Want a cookie?

OpenStudy (anonymous):

why not ??

OpenStudy (anonymous):

yes, i noticed it now. the fact is that i'm really sleepy and i have an exam of calc 1 today in a few hours (and studied nothing)...

OpenStudy (kainui):

Sure, ask me anything and I'll tell you anything you need to know and explain it in the best possible way I can so that hopefully it just "makes sense" and you don't have to memorize it.

OpenStudy (anonymous):

\[\int\limits_{}^{}e ^{x ^{3}}dx\] @Kainui

OpenStudy (kainui):

Lol go away kid.

OpenStudy (anonymous):

haha ... solve it ...!!

OpenStudy (anonymous):

x^3 is the u, no?

OpenStudy (kainui):

http://www.wolframalpha.com/input/?i=e%5Ex%5E3dx&t=crmtb01 There you go, now go away, I'm trying to actually help here not show off.

OpenStudy (kainui):

Don't read what he puts, it's actually a really hard integral similar to the Gaussian. It's not worth your time. @JGMach

OpenStudy (anonymous):

ya m sorry ...!!

OpenStudy (anonymous):

kainui, i was goig to ask you to solve a two integrals that i had a great doubt on how to solve them, but i took so long typing that i solved both in the middle of the process. the virtual keyboard is really a pellet to use. i need to buy a new one urgently! ps: sorry for my bad english

OpenStudy (anonymous):

i have another one here. how do i solve \[\int\limits \cos [\ln (x)]/(x.\sin [\ln(x)])dx\]?

OpenStudy (anonymous):

@Kainui